OpenStudy (anonymous):

Solve the equation below. Write out each step as you solve. 5. -6 (2x - 3) + 15 = 9 I can't get this! Please help! :-(

OpenStudy (anonymous):

First, you want to isolate the variable, so you'll want to get rid of the +15 on the left side by subtracting 15 from both sides. \[-6(2x-3) = 9 -15\] \[-6(2x-3) = -4\] Now, when you look at this equation, the -6 looks like it's attached to the variable x. How do you get cancel -6?

OpenStudy (anonymous):

@onemillion6, isn't 9-15 -6 not -4?

OpenStudy (anonymous):

@mipie13 haha you're correct! I'm sorry!

OpenStudy (anonymous):

So then I use the distributive property?

OpenStudy (anonymous):

You could! Or you can divide through by -6 because (with your correction) -6(2x-3) = -6. If you divide both sides by -6, they cancel!